OpenStudy (anonymous):

Which of the following is a pair of equivalent fractions? 3/4 and 8/12 2/4 and 10/25 6/9 and 12/14 1/8 and 2/4

OpenStudy (anonymous):

(E) none of the above

OpenStudy (anonymous):

b. 2/5 and 10/25 ps. i didn't know i hit he wrong number

OpenStudy (anonymous):

3/4 = 9/12 2/5 = 10/25 6/7 = 12/14 4/8 = 2/4

OpenStudy (anonymous):

PLEASE HELP ME!!

OpenStudy (anonymous):

can anybody help me?